OpenStudy (anonymous):

Find an equation of the line containing the given point and with the given slope. a.) (1,2), m=3 b) (0,-4), m=-2

OpenStudy (phi):

write down what you know, starting with y= mx+b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok y=mx+b y=3x+ would it be the 2 even though the given point is (1,2)

OpenStudy (phi):

pretty close. leave in the b y= 3x+b now "solve for b" using the (x,y) pair they gave you. that means the first step is to replace x and y with the numbers. and leave in the b

OpenStudy (anonymous):

y=mx+b y=3x+b 2=3(1)+b

OpenStudy (phi):

now simplify 3(1) 3*1 is 3 so you have 2= 3+b I would add -3 to both sides

OpenStudy (anonymous):

-1=b

OpenStudy (phi):

now you know m and b, so you can write down the equation of the line

OpenStudy (anonymous):

so it would be: 2=3x-1

OpenStudy (phi):

almost. try again

OpenStudy (anonymous):

2=3x+-1

OpenStudy (phi):

you start with the form y= mx+b as the equation you want. Using all the info they give you, you figure out m and b. Replace m and b with the numbers, and that is the equation of the line

OpenStudy (anonymous):

so y=3x+b from there i plug in my given points 2=3(1)+b from there i find b by adding -3 to both sides -1=b so my equation would be 2=3-1???????

OpenStudy (phi):

you found m=3 (actually, they told you that) and you figured out b= -1 if you go back to the bottom of the previous post, you asked how to find the equation of a line if you know the slope (m) and the y-intercept (b). Look at that. But 2=3-1??????? is not the equation of a line. It is a true equation 2 is 3-1, but there is no x and y. The equation of a line tells you how to find y when you are given x. you need a y and an x !

OpenStudy (phi):

If it helps, after you know m=3 and b= -1. Erase everything from your brain and start with y = mx+b what is the equation of the line when m= 3 and b=-1 ?

OpenStudy (anonymous):

y=3x-1?

OpenStudy (phi):

yes. before you had 2=3x-1 which was close. but the 2 should be a y notice that you started with y= mx+b and replace m with 3 (they told you that) to get y= 3x+b then to figure out b, you used the fact that (1,2) is on the line, so it must be true that 2= 3(1)+b and b must be, after solving, -1 next, you write down the important facts: m=3 and b= -1 and the form y= mx+b fill in for m and b and you get the answer y= 3x -1

OpenStudy (phi):

now try question (b)

OpenStudy (anonymous):

(0,4), m=-2 y=mx+b y=-2x+b 4=-2x(0) +b 4=0+b so the answer is y=-2x + 4

OpenStudy (anonymous):

that was supposed to say 4=-2(0) +b without the x in there

OpenStudy (phi):

yes. notice that point (0,4) has 4 on the y-axis. 4 is the y-intercept so you could just write y= -2x + 4. But doing the way you did also works.